Three friends Alice(A), Bob(B), and Charlie(C) are planning to buy a TV.

TV’s are available in all sizes (size is the diagonal length of the display in

inches). Each friend i has a private valuation of the form

vi(x) = aix − x^3/36

for a TV of size x, where i ∈ {A, B, C}

(a)

Use the VCG mechanism to decide which size of the TV should be

bought. Describe the size as a function of aA, aB, aC.

(b)

Compute the payments for each friend under the VCG mechanism in

the situation where Alice has aA = 10, Bob has aB = 40, and Charlie

has aC = 50.


1
Expert's answer
2021-11-17T15:40:12-0500

a)

VCG mechanism:

outcomes:

"x^*(v)=arg\ max_x(\sum v_i(x))"

"x^*(v_{-i})=arg\ max_x(\sum_{j\neq i} v_j(x))"

transfers: 

Agent i receives

"t_i(v)=\sum_{j\neq i}v_j(x^*(v))-\sum_{j\neq i}v_j(x^*(v_{-i}))"


for aA size of the TV:

x is root of equation "v_A(x) = a_Ax \u2212 x^3\/36"

for aB size of the TV:

x is root of equation "v_B(x) = a_Bx \u2212 x^3\/36"

for aC size of the TV:

x is root of equation "v_C(x) = a_Cx \u2212 x^3\/36"


b)

for aA = 10 :

"v_A(x) = 10x \u2212 x^3\/36"

for aB = 40 :

"v_B(x) = 40x \u2212 x^3\/36"


for aC = 50 :

"v_C(x) = 50x \u2212 x^3\/36"


"\sum v_i(x)=100x-x^3/12"


"(\sum v_i(x))'=100-x^2/4=0"

"x=20"

outcome that maximizes the sum of values:

"x^*(v)=20"


for Alice(A):

"\sum_{j\neq i} v_j(x)=90x-x^3/18"

"\sum_{j\neq i} v_j(x)=90-x^2/6=0"

"x^*(v_{-A})=\sqrt{540}=23.24"

Alice receives:

"t_A=90\cdot20-20^3/18-(90\cdot23.24-23.24^3/18)=-38.72"


for Bob(B):

"\sum_{j\neq i} v_j(x)=60x-x^3/18"

"\sum_{j\neq i} v_j(x)=60-x^2/6=0"

"x^*(v_{-B})=\sqrt{360}=18.97"

Bob receives:

"t_B=60\cdot20-20^3/18-(60\cdot18.97-18.97^3/18)=-3.39"


for Charlie(C):

"\sum_{j\neq i} v_j(x)=50x-x^3/18"

"\sum_{j\neq i} v_j(x)=50-x^2/6=0"

"x^*(v_{-A})=\sqrt{300}=17.32"

Charlie receives:

"t_C=50\cdot20-20^3/18-(50\cdot17.32-17.32^3/18)=-21.79"


So, Alice has to pay 38.72, Bob has to pay 3.39, Charlie has to pay 21.79